# Pyramidal Flow Matching for Efficient Video Generative Modeling

### v1

@@ Abstract
Video generation requires modeling a vast spatiotemporal space, which demands significant computational resources and data usage. To reduce the complexity, the prevailing approaches employ a cascaded architecture to avoid direct training with full resolution. Despite reducing computational demands, the separate optimization of each sub-stage hinders knowledge sharing and sacrifices flexibility. This work introduces a unified pyramidal flow matching algorithm. It reinterprets the original denoising trajectory as a series of pyramid stages, where only the final stage operates at the full resolution, thereby enabling more efficient video generative modeling. Through our sophisticated design, the flows of different pyramid stages can be interlinked to maintain continuity. Moreover, we craft autoregressive video generation with a temporal pyramid to compress the full-resolution history. The entire framework can be optimized in an end-to-end manner and with a single unified Diffusion Transformer (DiT). Extensive experiments demonstrate that our method supports generating high-quality 5-second (up to 10-second) videos at 768p resolution and 24 FPS within 20.7k A100 GPU training hours. All code and models will be open-sourced at https://pyramid-flow.github.io.
@@ 1 Introduction
Based on the aforementioned pyramidal representations, we introduce a novel pyramidal flow matching algorithm that builds upon recent prevalent flow matching framework (Lipman et al., 2023; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023). Specifically, we devise a piecewise flow for each pyramid resolution, which together form a generative process from noise to data. The flow within each pyramid stage takes a similar formulation, interpolating between a pixelated (compressed) and noisier latent and a pixelate-free (decompressed) and cleaner latent. Thanks to our sophisticated design, they can be jointly optimized by the unified flow matching objective in a single Diffusion Transformer (DiT) (Peebles & Xie, 2023), allowing simultaneous generation and decompression of visual content without multiple separate models. During inference, the output of each stage is renoised by a corrective Gaussian noise, which contributes to maintaining the continuity of the probability path between successive pyramid stages. Furthermore, we formulate the video generation in an autoregressive manner that iteratively predicts the next video latent conditioned on previously generated history. Given the high redundancy in the full-resolution history, we curate a temporal pyramid sequence using progressively compressed, lower-resolution history as conditions, thereby further reducing the token count and improving training efficiency.
@@ 2 Related Work
Image Pyramids (Adelson et al., 1984) have been studied extensively in visual representation learning (Lowe, 2004; Dalal & Triggs, 2005; Lin et al., 2017; Wang et al., 2020). For generative models, the idea is explored by cascaded diffusion models that first generate at low resolution and then perform super-resolution (Ho et al., 2022b; Saharia et al., 2022; Pernias et al., 2024; Teng et al., 2024). It is also extended to video by spatiotemporal super-resolution (Ho et al., 2022a; Singer et al., 2023). However, they require training several separate models, which prevents knowledge sharing. Possible unified modeling solutions for pyramids include hierarchical architectures (Rombach et al., 2022; Crowson et al., 2024; Hatamizadeh et al., 2024) or via next-token prediction (Yu et al., 2023; Tian et al., 2024), but involve architectural changes. Instead, we propose a simple flow matching objective that allows joint training of pyramids, thus facilitating efficient video generative modeling.
@@ 3.1 Preliminaries on Flow Matching
and . Notably, flow matching can be flexibly extended to interpolate between distributions other than standard Gaussians. This enables us to devise a sophisticated flow matching algorithm that specializes in reducing the computational cost of video generative modeling.
@@ 3.2 Pyramidal Flow Matching
The main challenge in video generative modeling is the spatio-temporal complexity, and we address its spatial complexity first. According to previous key observation in Fig. 1, the initial generation steps are usually very noisy and less informative, and thus may not need to operate at full resolution. This motivates us to study a spatially compressed pyramidal flow, illustrated in Fig. 2.
@@ 4.2 Efficiency
| Open-Sora Plan v1.1 | ✓ | 78.00 | 80.91 | 66.38 | 98.28 | 47.72 | ||||||||||||
@@ 4.3 Main Results
Text-to-Video Generation. We first evaluate the text-to-video generation capability of the proposed method. For each text prompt, a 5-second 121 frames video is generated for evaluation. The detailed quantitative results on VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024) are summarized in Tables 1 and 2, respectively. Overall, our method surpasses all the compared open-sourced video generation baselines in these two benchmarks. Even with only publicly accessible video data in training, it achieves comparable performance to commercial competitors trained on much larger proprietary data like Kling (Kuaishou, 2024) and Gen-3 Alpha (Runway, 2024). In particular, we demonstrated exceptional performance in quality score (84.74 vs. 84.11 of Gen-3), and motion smoothness in VBench, which are crucial criteria in reflecting the visual quality of generated videos. When evaluated in EvalCrafter, our method achieves better visual and motion quality scores than most compared methods. The semantic score is relatively lower than others, mainly because we use coarse-grained synthetic captions,
which can be improved with more accurate video captioning. We also present some generated 5–10 second videos in Fig. 5, showing cinematic visual quality and validate the efficacy of pyramidal flow matching. More visualizations are provided in Section C.3.
User study. While quantitative evaluation scores reflect the video generation capability to some extent, they may not align with human preferences for visual quality. Hence, an additional user study is conducted to compare our performance with six baseline models, including CogVideoX (Yang et al., 2024) and Kling (Kuaishou, 2024). We utilized 50 prompts sampled from VBench and asked 20+ participants to rank each model according to the aesthetic quality, motion smoothness, and semantic alignment of the generated videos. As seen in Fig. 4, our method is preferred over open-source models such as Open-Sora and CogVideoX-2B especially in terms of motion smoothness. This is due to the substantial token savings achieved by pyramidal flow matching, enabling generation of 5-second (up to 10-second) 768p videos at 24 fps, while the baselines usually support video synthesis of similar length only at 8 fps. The detailed user study settings are presented in Appendix B.
@@ 5 Conclusion
This work presents an efficient video generative modeling framework based on pyramidal visual representations. In contrast to cascaded diffusion models that use separate models for different image pyramids to improve efficiency, we propose a unified pyramidal flow matching objective that simultaneously generates and decompresses visual content across pyramid stages with a single model, effectively facilitating knowledge sharing. Furthermore, a temporal pyramid design is introduced to reduce computational redundancy in the full-resolution history of a video. The proposed method is extensively evaluated on VBench and EvalCrafter, demonstrating its advantageous performance. All code and model weights will be open-sourced at https://pyramid-flow.github.io.
Acknowledgements. We thank Jinghan Li for assisting with the user study.
@@ Appendix A Derivation
This section provides detailed derivation for Eq. 15 that handles jump points in the spatial pyramid.
@@ Appendix B Experimental Settings
Baseline Methods. For VBench (Huang et al., 2024), we compare with eight baseline methods, including Open-Sora Plan V1.1 (PKU-Yuan Lab et al., 2024), Open-Sora 1.2 (Zheng et al., 2024), VideoCrafter2 (Chen et al., 2024b), Gen-2 (Runway, 2023), Pika 1.0 (Pika, 2023), T2V-Turbo (Li et al., 2024), CogVideoX (Yang et al., 2024), Kling (Kuaishou, 2024), and Gen-3 Alpha (Runway, 2024). Among them, Open-Sora Plan, Open-Sora, CogVideo-X, Kling and Gen-3 Alpha can generate long videos. For EvalCrafter (Liu et al., 2024), our model is compared to six baselines, including ModelScope (Wang et al., 2023a), Show-1 (Zhang et al., 2024), LaVie (Wang et al., 2023b), VideoCrafter2 (Chen et al., 2024b), Pika 1.0 (Pika, 2023), and Gen-2 (Runway, 2023). The above models are all based on full-sequence diffusion, while our method combines the merits of autoregressive generation and flow generative models to achieve better training efficiency of video generation.
User Study. To complement the quantitative evaluation in the main paper, we conduct a rigorous user study to collect human preferences for these generative models. To accomplish this, we sample 50 prompts from the VBench prompt list and randomly sample one generated video for each prompt from the baseline model. In total, six baseline models are considered, including Open-Sora Plan V1.1 (PKU-Yuan Lab et al., 2024), Open-Sora 1.2 (Zheng et al., 2024), Pika 1.0 (Pika, 2023), CogVideoX-2B and 5B (Yang et al., 2024), and Kling (Kuaishou, 2024). We then pair these results with our generated video and ask the participant to rank their preference among three dimensions: aesthetic quality, motion smoothness, and semantic alignment, each of which represents a crucial aspect of video quality. The interface for the user study is exemplified in Fig. 9, where the user accepts a prompt and two generated videos (with the unnecessary information cropped, such as a watermark indicating which model it belongs to), and chooses between which model is better in the three dimensions. We distribute the user study to more than 20 participants, and collect a total of 1411 valid preference choices, ensuring its effectiveness. The results of this user study are presented in Fig. 4, where our model shows a very competitive performance among the compared baselines.
@@ C.1 Quantitative Results
| Open-Sora Plan v1.1 | 95.73 | 96.73 | 99.03 | 98.28 | 47.72 | 56.85 | 62.28 | 76.30 | ||||||||||||||||
@@
| Open-Sora Plan v1.1 | 40.35 | 86.8 | 89.19 | 53.11 | 27.17 | 22.90 | 23.87 | 26.52 | ||||||||||||||||
@@
VBench (Huang et al., 2024). The full experimental results on VBench are shown in Table 4. As can be observed, our model achieves leading or highly competitive results among open-source and commercial competitors, especially for the metrics related to motion quality. For example, the dynamic degree metric of our model ranks 2nd among all models at 64.63, validating the effectiveness of our generative model in learning temporal dynamics. For the rest of the metrics, our results are also generally superior to the open-source Open-Sora Plan v1.1 (PKU-Yuan Lab et al., 2024) and Open-Sora 1.2 (Zheng et al., 2024), with significantly lower training computational cost as mentioned earlier. We also note that half of our results even outperformed the recent CogVideoX-5B (Yang et al., 2024), which is based on a larger DiT model, demonstrating its modeling capacity. On the other hand, our model performs relatively inferior on metrics such as color and appearance style, which is more related to the image generation capabilities and finer-grained prompt following. This is largely due to our video captioning procedure based on video LLMs which tends to produce coarse-grained captions, thus dampening these abilities. Nevertheless, thanks to our autoregressive generation framework, which decomposes video generation into first frame generation and subsequent frame generation, these image quality issues can be addressed separately with additional well-captioned image data in future training stages. Similarly, due to the SD3-Medium weight initialization, which is infamous for its human structure, our method achieves a relatively low score in human action, which could be addressed by switching to other base models or training from scratch.
EvalCrafter (Liu et al., 2024). The raw metrics on EvalCrafter are provided in Table 5. Overall, our model delivers highly competitive performance on the majority of metrics, outperforming many previous open-source and closed-source models. In particular, the motion AC score of our method which is relevant to the temporal motion quality ranks 2nd among all methods, justifying the capacity of our pyramid designs to learn complex spatiotemporal patterns in video. Our method also demonstrates superiority over several other metrics related to semantic alignment, including BLIP-BLUE and CLIP score. Placing top two in both metrics among the models compared, including the closed-source Gen-2 (Runway, 2023), confirms the advantages of our model in text-to-video semantic alignment. The only metric where our model performs poorly is face consistency, which is due to the temporal pyramid design adopted for compressing the history condition. We view this as an issue that can potentially be addressed by more sophisticated temporal compression schemes.

### v2

@@ Abstract
Video generation requires modeling a vast spatiotemporal space, which demands significant computational resources and data usage. To reduce the complexity, the prevailing approaches employ a cascaded architecture to avoid direct training with full resolution latent. Despite reducing computational demands, the separate optimization of each sub-stage hinders knowledge sharing and sacrifices flexibility. This work introduces a unified pyramidal flow matching algorithm. It reinterprets the original denoising trajectory as a series of pyramid stages, where only the final stage operates at the full resolution, thereby enabling more efficient video generative modeling. Through our sophisticated design, the flows of different pyramid stages can be interlinked to maintain continuity. Moreover, we craft autoregressive video generation with a temporal pyramid to compress the full-resolution history. The entire framework can be optimized in an end-to-end manner and with a single unified Diffusion Transformer (DiT). Extensive experiments demonstrate that our method supports generating high-quality 5-second (up to 10-second) videos at 768p resolution and 24 FPS within 20.7k A100 GPU training hours. All code and models are open-sourced at https://pyramid-flow.github.io.
@@ 1 Introduction
Based on the aforementioned pyramidal representations, we introduce a novel pyramidal flow matching algorithm that builds upon recent prevalent flow matching framework (Lipman et al., 2023; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023). Specifically, we devise a piecewise flow for each pyramid resolution, which together form a generative process from noise to data. The flow within each pyramid stage takes a similar formulation, interpolating between a pixelated (compressed) and noisier latent and a pixelate-free (decompressed) and cleaner latent. Through our design, they can be jointly optimized by the unified flow matching objective in a single Diffusion Transformer (DiT) (Peebles & Xie, 2023), allowing simultaneous generation and decompression of visual content without multiple separate models. During inference, the output of each stage is renoised by a corrective Gaussian noise that maintains the continuity of the probability path between stages. Furthermore, we formulate the video generation in an autoregressive manner, iteratively predicting the next video latent conditioned on the generated history. Given the high redundancy in the full-resolution history, we curate a temporal pyramid sequence using progressively compressed, lower-resolution history as conditions, further reducing the number of tokens and improving training efficiency.
@@ 2 Related Work
Image Pyramids (Adelson et al., 1984) have been studied extensively in visual representation learning (Lowe, 2004; Dalal & Triggs, 2005; Lin et al., 2017; Wang et al., 2020). For generative models, the idea is explored by cascaded diffusion models that first generate at low resolution and then perform super-resolution (Ho et al., 2022b; Saharia et al., 2022; Zhang et al., 2023b; Gu et al., 2023; Pernias et al., 2024; Teng et al., 2024), and later extended to video (Ho et al., 2022a; Singer et al., 2023). However, they require training several separate models, which prevents knowledge sharing. Possible unified modeling solutions for pyramids include hierarchical architectures (Rombach et al., 2022; Crowson et al., 2024; Hatamizadeh et al., 2024) or via next-token prediction (Yu et al., 2023; Tian et al., 2024), but involve architectural changes. Instead, we propose a simple flow matching objective that allows joint training of pyramids, thus facilitating efficient video generative modeling.
@@ 3.1 Preliminaries on Flow Matching
and . Notably, flow matching can be flexibly extended to interpolate between distributions other than standard Gaussians. This enables us to devise a new flow matching algorithm that specializes in reducing the computational cost of video generative modeling.
@@ 3.2 Pyramidal Flow Matching
The main challenge in video generative modeling is the spatio-temporal complexity, and we address its spatial complexity first. According to previous key observation in Fig. 1, the initial generation steps are usually very noisy and less informative, and thus may not need to operate at full resolution latent. This motivates us to study a spatially compressed pyramidal flow, illustrated in Fig. 2.
@@ 4.2 Efficiency
| Open-Sora Plan v1.3 | ✓ | 77.23 | 80.14 | 65.62 | 99.05 | 30.28 | ||||||||||||
@@ 4.3 Main Results
Text-to-Video Generation.
We first evaluate the text-to-video generation capability of the proposed method. For each text prompt, a 5-second 121 frames video is generated for evaluation. The detailed quantitative results on VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024) are summarized in Tables 1 and 2, respectively. Overall, our method surpasses all the compared open-sourced video generation baselines in these two benchmarks. Even with only publicly accessible video data in training, it achieves comparable performance to commercial competitors trained on much larger proprietary data like Kling (Kuaishou, 2024) and Gen-3 Alpha (Runway, 2024). In particular, we demonstrated exceptional performance in quality score (84.74 vs. 84.11 of Gen-3), and motion smoothness in VBench, which are crucial criteria in reflecting the visual quality of generated videos. When evaluated in EvalCrafter, our method achieves better visual and motion quality scores than most compared methods. The semantic score is relatively lower than others, mainly because we use coarse-grained synthetic captions, which can be improved with more accurate video captioning. We also present some generated 5–10 second videos in Fig. 5, showing cinematic visual quality and validate the efficacy of pyramidal flow matching. More visualizations are provided in Section C.3.
User study. While quantitative evaluation scores reflect the video generation capability to some extent, they may not align with human preferences for visual quality. Hence, an additional user study is conducted to compare our performance with six baseline models, including CogVideoX (Yang et al., 2024) and Kling (Kuaishou, 2024). We utilized 50 prompts sampled from VBench and asked 20+ participants to rank each model according to the aesthetic quality, motion smoothness, and semantic alignment of the generated videos. As seen in Section 4.3, our method is preferred over open-source models such as Open-Sora and CogVideoX-2B especially in terms of motion smoothness. This is due to the substantial token savings achieved by pyramidal flow matching, enabling generation of 5-second (up to 10-second) 768p videos at 24 fps, while the baselines usually support video synthesis of similar length only at 8 fps. The detailed user study settings are presented in Appendix B.
@@ 5 Conclusion
This work presents an efficient video generative modeling framework based on pyramidal visual representations. In contrast to cascaded diffusion models that use separate models for different image pyramids to improve efficiency, we propose a unified pyramidal flow matching objective that simultaneously generates and decompresses visual content across pyramid stages with a single model, effectively facilitating knowledge sharing. Furthermore, a temporal pyramid design is introduced to reduce computational redundancy in the full-resolution history of a video. The proposed method is extensively evaluated on VBench and EvalCrafter, demonstrating advantageous performance.
Reproducibility Statement. Our code and models are open-sourced at https://pyramid-flow.github.io. The experimental settings are detailed in Sections 4.1 and B.
Acknowledgments
The work was supported by National Key R&D Program of China (2022ZD0160300), an internal grant of Peking University (2024JK28), a grant from Kuaishou (No. DJHL-20240809-115) and NSF China (No. 62276004).
@@ Appendix A Derivation
This section provides a detailed derivation for Eq. 15 that handles jump points in the spatial pyramid. For quick lookup,Table 3 summarizes the used notations.
| Symbol | Description |
| Data latent at full resolution | |
| Noise at full resolution | |
| Number of pyramid stages | |
| Timestep at endpoint of -th pyramid stage | |
| Timestep at starting point of -th pyramid stage | |
| Noisy latent at endpoint of -th stage | |
| Noisy latent at starting point of -th stage | |
| Noisy latent at timestep | |
| Noise at the resolution of current stage | |
| Upsampling function, e.g. nearest-neighbor | |
| Downsampling function, e.g. bilinear |
@@ Appendix B Experimental Settings
Baseline Methods. For VBench (Huang et al., 2024), we compare with eight baseline methods, including Open-Sora Plan V1.3 (PKU-Yuan Lab et al., 2024), Open-Sora 1.2 (Zheng et al., 2024), VideoCrafter2 (Chen et al., 2024b), Gen-2 (Runway, 2023), Pika 1.0 (Pika, 2023), T2V-Turbo (Li et al., 2024), CogVideoX (Yang et al., 2024), Kling (Kuaishou, 2024), and Gen-3 Alpha (Runway, 2024). Among them, Open-Sora Plan, Open-Sora, CogVideo-X, Kling and Gen-3 Alpha can generate long videos. For EvalCrafter (Liu et al., 2024), our model is compared to six baselines, including ModelScope (Wang et al., 2023a), Show-1 (Zhang et al., 2023a), LaVie (Wang et al., 2023b), VideoCrafter2 (Chen et al., 2024b), Pika 1.0 (Pika, 2023), and Gen-2 (Runway, 2023). The above models are all based on full-sequence diffusion, while our method combines the merits of autoregressive generation and flow generative models to achieve better training efficiency of video generation.
User Study. To complement the quantitative evaluation in the main paper, we conduct a rigorous user study to collect human preferences for these generative models. To accomplish this, we sample 50 prompts from the VBench prompt list and randomly sample one generated video for each prompt from the baseline model. In total, six baseline models are considered, including Open-Sora Plan V1.1 (PKU-Yuan Lab et al., 2024), Open-Sora 1.2 (Zheng et al., 2024), Pika 1.0 (Pika, 2023), CogVideoX-2B and 5B (Yang et al., 2024), and Kling (Kuaishou, 2024). We then pair these results with our generated video and ask the participant to rank their preference among three dimensions: aesthetic quality, motion smoothness, and semantic alignment, each of which represents a crucial aspect of video quality. The interface for the user study is exemplified in Fig. 9, where the user accepts a prompt and two generated videos (with the unnecessary information cropped, such as a watermark indicating which model it belongs to), and chooses between which model is better in the three dimensions. We distribute the user study to more than 20 participants, and collect a total of 1411 valid preference choices, ensuring its effectiveness. The results of this user study are presented in Section 4.3, where our model shows a very competitive performance among the compared baselines.
@@ C.1 Quantitative Results
| Open-Sora Plan v1.3 | 97.79 | 97.24 | 99.20 | 99.05 | 30.28 | 60.42 | 56.21 | 85.56 | ||||||||||||||||
@@
| Open-Sora Plan v1.3 | 43.58 | 86.8 | 79.30 | 51.61 | 36.73 | 20.03 | 22.47 | 24.47 | ||||||||||||||||
@@
VBench (Huang et al., 2024). The full experimental results on VBench are shown in Table 5. As can be observed, our model achieves leading or highly competitive results among open-source and commercial competitors, especially for the metrics related to motion quality. For example, the dynamic degree metric of our model ranks 2nd among all models at 64.63, validating the effectiveness of our generative model in learning temporal dynamics. For the rest of the metrics, our results are also generally superior to the open-source Open-Sora Plan v1.3 (PKU-Yuan Lab et al., 2024) and Open-Sora 1.2 (Zheng et al., 2024), with significantly lower training computational cost as mentioned earlier. We also note that half of our results even outperformed the recent CogVideoX-5B (Yang et al., 2024), which is based on a larger DiT model, demonstrating its modeling capacity. On the other hand, our model performs relatively inferior on metrics such as color and appearance style, which is more related to the image generation capabilities and finer-grained prompt following. This is largely due to our video captioning procedure based on video LLMs which tends to produce coarse-grained captions, thus dampening these abilities. Nevertheless, thanks to our autoregressive generation framework, which decomposes video generation into first frame generation and subsequent frame generation, these image quality issues can be addressed separately with additional well-captioned image data in future training stages. Similarly, due to the SD3-Medium weight initialization, which is infamous for its human structure, our method achieves a relatively low score in human action, which could be addressed by switching to other base models or training from scratch.
EvalCrafter (Liu et al., 2024). The raw metrics on EvalCrafter are provided in Table 6. Overall, our model delivers highly competitive performance on the majority of metrics, outperforming many previous open-source and closed-source models. In particular, the motion AC score of our method which is relevant to the temporal motion quality ranks 2nd among all methods, justifying the capacity of our pyramid designs to learn complex spatiotemporal patterns in video. Our method also demonstrates superiority over several other metrics related to semantic alignment, including BLIP-BLUE and CLIP score. Placing top two in both metrics among the models compared, including the closed-source Gen-2 (Runway, 2023), confirms the advantages of our model in text-to-video semantic alignment. The only metric where our model performs poorly is face consistency, which is due to the temporal pyramid design adopted for compressing the history condition. We view this as an issue that can potentially be addressed by better temporal compression schemes.
@@ C.3 Visualization
This section presents additional qualitative results for our text-to-video generation in comparison to the recent leading models including Gen-3 Alpha (Runway, 2024), Kling (Kuaishou, 2024) and CogVideoX (Yang et al., 2024). The uniformly sampled frames from the generated videos are shown in Figs. 15 and 14, in which our videos are generated at 5s, 768p, 24fps. Overall, we observe that despite being trained only on publicly available data and using a small computational budget, our model yields a highly competitive visual aesthetics and motion quality among the baselines.
Specifically, the results highlight the following characteristics of our model: (1) Through generative pre-training, our model is capable of generating videos of cinematic quality and reasonable content. For example, in Fig. 14(a), our generative video shows a mushroom cloud resulting from “a massive explosion” taking place in “the surface of the earth”, creating a sci-fi movie atmosphere. However, the current model is not fully faithful to some prompts such as the “salt desert” in Fig. 14(b), which could be addressed by curating more high-quality caption data. (2) Despite that our model has only 2B parameters initialized from SD3-Medium (Esser et al., 2024), it clearly outperforms CogVideoX-2B of the same model size with additional training data, and is even comparable to the 5B full version in some aspects. For example, in Figs. 15(a) and 15(b), only our model and its 5B version are capable of generating reasonable sea waves according to the input prompt, while its 2B variant merely illustrates an almost static sea surface. This is largely attributed to our proposed pyramidal flow matching in improving training efficiency. Overall, these results validate the effectiveness of our approach in modeling complex spatiotemporal patterns through the spatial and temporal pyramid designs. Our generated videos are best-viewed at https://pyramid-flow.github.io. Since the autoregressive video generation model natively generates a high-quality image as the first frame, pyramid-flow can also be applied to text-to-image generation. Even with only a few million training images, it can show excellent visual quality, see Fig. 12(a) for the generated images.
C.4 Toy Experiment of Coupling Noise
To validate the effectiveness of coupled sampling in Eqs. 9 and 10, we illustrate two variants of piecewise flow matching in a toy experiment that considers mapping a few data points to uniform distribution. Two different coupling designs are considered within each time window, namely our coupling vs. random coupling. It can be seen that our coupled sampling strategy produces much more straight flow trajectories.
The rationale for improving straightness by coupling noise is that: the straightness of the flow trajectory is usually compromised when there are intersections. Sampling the endpoints independently (as in vanilla flow matching) creates random directions for each trajectory and leads to intersections. Instead, by coupling the sampling of these endpoints as in Eq. 9 and Eq. 10, we can create more organized, possibly parallel trajectories with fewer intersections, thus improving straightness. As illustrated in Fig. 13, where coupling noise indeed leads to more straight flow trajectories.
Appendix D Limitations
Our method only supports autoregressive generation and cannot be extended to keyframe interpolation or video interpolation. In addition, we noticed that the temporal pyramid designs to improve training efficiency can sometimes lead to subtle subject inconsistency, especially over the long term. While this is not a prevalent problem, we believe that developing better temporal compression methods is critical to the broader applicability of autoregressive video generative model. In addition, improving inference efficiency towards real-time is an intriguing problem (Yin et al., 2024).